OpenStudy (anonymous):

solve the system by substitution 2x+y=-11, 3x-4y=11 a. (3,5) b. (-5,-3) c. (-3,-5) d. (5,3)

OpenStudy (anonymous):

isolate y in the first equation

OpenStudy (anonymous):

u get y = -2x-11, which u then substitute into the second equation

OpenStudy (anonymous):

the derived second equation: 3x - 4(-2x-11) = 11, do ur algebra, solve for x then plug it back in any of the two equation to get the y value, after that ur all done

OpenStudy (anonymous):

is it 15 so would it be d @11calcBC

OpenStudy (anonymous):

wuts ur x value? from my equation

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

3x - 4(-2x-11) = 11, after distributing u get 3x + 8x + 44 = 11, simplify, 11x = -33, wuts x?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yes there we go

OpenStudy (anonymous):

u understand now?

OpenStudy (anonymous):

yep so its would be c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks